OpenStudy (anonymous):

Write the equation for the line which passes through the point (4, 1) and is perpendicular to the graph of y = x – 3. Write the equation in slope-intercept form. No idea how to do this. Help please! :(

OpenStudy (anonymous):

you need a point and a slope the point you are given, it is (4,1) the slope you find the line \(y=x-3\) has slope 1, and so the perpendicular line will have slope \(-1\) because it is the negative reciprocal

OpenStudy (anonymous):

then you need the point slope formula \(y-y_1=m(x-x_1)\) with \[m=-1,x_1=4,y_x=1\] make a direct substitution to get your answer